# Exploring Selective Layer Fine-Tuning in Federated Learning

### v1

@@ 6 Conclusions
Appendix A Algorithm Implementation
The training process of the proposed selective layer fine-tuning framework for FL is summarized in Algorithm 1.
Algorithm 1 Selective Layer Fine-tuning in FL
  Input: The pre-trained global model $\theta^{0}$
  for $t=0,1,\dots,T-1$ do
     Sample a set of clients $\mathcal{S}^{t}$;
     for each client $i$ in $\mathcal{S}^{t}$ do
[…]
     end for
[…]
  end for
  Return: The global model $\theta^{T}$
Appendix B Convergence Analysis: Full Proofs
B.1 Useful Lemmas
B.1.1 One-Round Loss Decay
Lemma B.1.
@@ B.1.1 One-Round Loss Decay
B.1.2 Analyzing $\mathcal{E}_{t}$: Proof of Lemma 4.6
@@
B.2 Convergence Analysis for Single-Step Case
@@
Using the result in Lemma B.1, we have:
@@
B.3 Convergence Analysis for Multi-Step Case
@@
In Lemma B.1, the term $\mathcal{T}_{1}$ is related to client drift caused by multiple local SGD steps, which can be upper bounded as follows:
@@
Following Lemma 22 in [39], $\mathcal{T}_{4}$ can be upper bounded as:
@@
$\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|^{2}\right]\leq\frac{4}{\eta\tau C^{\prime}}\left[\mathbb{E}[f(\theta^{t})]-\mathbb{E}[f(\theta^{t+1})]\right]+(\frac{1}{\eta\tau\gamma C^{\prime}}+2)\mathcal{E}_{t}+\frac{4A_{\tau}}{C^{\prime}}\sigma^{2}.$ (66)
@@
Appendix C Discussion on Computational and Communication Costs
Computational cost
Since all methods require the same forward operations, we focus on comparing the computational cost of backward operations in different methods. Consider that each client selects a layer to fine-tune a model with $L$ layers. For simplicity, we assume each layer requires $b$ FLOPs of backward operations. The average computational cost of the proposed layer selection method is calculated as:
$\text{Cost}_{\text{sel}}=\underbrace{b(L-1)}_{\text{Select}}+\underbrace{b\tau}_{\text{Fine-tune}}=b(\tau+L-1),$ (70)
$\text{Cost}_{\text{full}}=bL\tau=\frac{L\tau}{\tau+L-1}\text{Cost}_{\text{sel}}.$ (71)
Meanwhile, the selection step of our method introduces slightly additional costs of $\frac{L-1}{\tau L}\text{Cost}_{\text{full}}$, which can be further reduced by evaluating the model on a smaller volume of data or making the selection decision at a lower frequency.
Communication time
The communication time is determined by the transmitted bits of different methods. The proposed method only transmits the selected layers, which is much less than uploading the full model. For example, assuming that different layers have the same number of parameters, the communication cost of the proposed method is $\frac{1}{L}$ of full model fine-tuning.
Cost of fine-tuning specific layers
When a layer is selected for fine-tuning, the client performs the forward operation through the whole model to obtain the loss and then conducts the backward operation to compute the gradient with respect to the selected layer. The gradient has the same shape as the weights of the selected layer, and thus the required FLOPs are proportional to the number of parameters in this specific layer. Since all methods require the same forward operations, we focus on comparing the computational cost of backward operations in different methods.
Empirical observation
The following table compares the computational cost (in TFLOPs) and communication cost (in transmitted MBits) of the proposed method with the full model fine-tuning. The numbers in brackets represent the cost of the layer selection strategy. The results show that the proposed method significantly reduces the computational cost and communication costs. Besides, the layer selection strategy takes as low as 3.4% of the computational cost.
Computational cost (TFLOPs) Ratio Transmission (MBits) Ratio
Full Model Fine-tuning 8.47 100% 2,811 100%
Proposed Method 2.24 (1.51) 26% (17%) 234 8.33%
Proposed (Sel. Period=2) 1.46 (0.75) 17% (9.5%) 234 8.33%
Proposed (Sel. Batch=1) 0.99 (0.30) 12% (3.4%) 234 8.33%
Summarization
To summarize, the computational cost and communication cost/time of the proposed method are much lower than the full model fine-tuning. Besides, the empirical observations demonstrate that the proposed method significantly reduces the computational and communication costs.
Appendix D Experimental Details
@@ D.1 Training Tasks
  • CIFAR-10 [47]: In this training task, we consider the label-skewed case where $P(y_{i})$ is different among clients. Following previous works [26], we adopt Dirichlet allocation with concentration parameter $\alpha=0.1$ among clients.
  • DomainNet [48]: DomainNet contains six domains of data samples, including clipart, real, sketch, infograph, painting, and quickdraw. In this training task, we consider the varying feature case where $P(x_{i}|y_{i})$ is different among clients. Following [61], each client is allocated random samples from only one domain.
On both tasks, we fine-tune a CLIP Vision Transformer (CLIP) model [5].
@@
  • XGLUE-NC [49]: This is a news classification task consisting of 10 classes. The news texts comprise five languages (English, Spanish, French, German, and Russian). We allocate one random language to each client, which naturally introduces domain shift among clients.
@@
  • QA: The QA datasets consist of four commonly used question-answering datasets, i.e., SCIQ [50], OpenbookQA [51], PIQA [52], ARC-Easy and ARC-Challenge [53]. The datasets are transformed into classification tasks where the model determines the correct answer for each question and corresponding choices. We allocate the samples from one random dataset to each client, indicating the domain shift among clients.
D.2 Implementation Details
The CLIP model is pre-trained on the DataComp dataset and is adapted from https://github.com/openai/CLIP; the XLM-Roberta-Base model is adapted from https://huggingface.co/xlm-roberta-base; the LLaMA-2-7B model is adapted from https://huggingface.co/meta-llama/LLaMA-7b-chat-hf. In all training tasks, we freeze the embedding layers of the model and fix the classifier as commonly selected layers [32]. The values of adopted hyperparameters are summarized in Table 4.
@@ D.2 Implementation Details
Appendix E Limitations
One main limitation of our work is the theoretical results in Theorem 4.7 are restricted to the case where Assumptions 4.1-4.3 hold. Besides, the proposed layer selection strategy introduces additional computation costs, as discussed in Appendix C.

### v2

@@ 5.3 Further Discussions
Comparisons regarding computational and communication costs   We compare the computational cost (in TFLOPs) and communication cost (in transmitted MBits) of the proposed method with the full model fine-tuning when adopting the CLIP model on the CIFAR-10. For the proposed method, we consider fine-tuning only one layer, as it is sufficient to achieve comparable accuracy with full model fine-tuning according to Table 1. The results in Table 3 evidence a substantial decrease in both computational and communication requirements when utilizing the proposed method. Besides, the layer selection strategy takes as low as 3.4% of the computational cost. Therefore, the aforementioned empirical observations demonstrate that the proposed method significantly reduces both the computational and communication costs.
Table 3: Comparisons of computational and communication costs when fine-tuning the CLIP model on the CIFAR-10 ($R=1$). The numbers in brackets represent the cost of the layer selection strategy.
Computational cost (TFLOPs) Ratio Transmission (MBits) Ratio
Full Model Fine-tuning 8.47 100% 2,811 100%
Proposed Method 2.24 (1.51) 26% (17%) 234 8.33%
Proposed (Sel. Period=2) 1.46 (0.75) 17% (9.5%) 234 8.33%
Proposed (Sel. Batch=1) 0.99 (0.30) 12% (3.4%) 234 8.33%
@@ 6 Conclusions
Appendix A Convergence Analysis: Full Proofs
A.1 Useful Lemmas
A.1.1 One-Round Loss Decay
Lemma A.1.
@@ A.1.1 One-Round Loss Decay
A.1.2 Analyzing $\mathcal{E}_{t}$: Proof of Lemma 4.6
@@
A.2 Convergence Analysis for Single-Step Case
@@
Using the result in Lemma A.1, we have:
@@
A.3 Convergence Analysis for Multi-Step Case
@@
In Lemma A.1, the term $\mathcal{T}_{1}$ is related to client drift caused by multiple local SGD steps, which can be upper bounded as follows:
@@
Following Lemma 22 in (Pillutla et al., 2022), $\mathcal{T}_{4}$ can be upper bounded as:
@@
$\mathbb{E}\left[\left\|\nabla f(\theta^{t})\right\|^{2}\right]\leq\frac{4}{\eta\tau C^{\prime}}\left[\mathbb{E}[f(\theta^{t})]-\mathbb{E}[f(\theta^{t+1})]\right]+\left(\frac{1}{\eta\tau\gamma C^{\prime}}+2\right)\mathcal{E}_{t}+\frac{4A_{\tau}}{C^{\prime}}\sigma^{2}.$ (68)
@@
Appendix B Experimental Details
@@ B.1 Training Tasks
  • CIFAR-10 (Krizhevsky & Hinton, 2009): In this training task, we consider the label-skewed case where $P(y_{i})$ is different among clients. Following previous works (Li et al., 2022), we adopt Dirichlet allocation with concentration parameter $\alpha=0.1$ among clients.
  • DomainNet (Peng et al., 2019): DomainNet contains six domains of data samples, including clipart, real, sketch, infograph, painting, and quickdraw. In this training task, we consider the varying feature case where $P(x_{i}|y_{i})$ is different among clients. Following (Li et al., 2021), each client is allocated random samples from only one domain.
On both tasks, we fine-tune a CLIP Vision Transformer (CLIP) model (Radford et al., 2021).
@@
  • XGLUE-NC (Liang et al., 2020): This is a news classification task consisting of 10 classes. The news texts comprise five languages (English, Spanish, French, German, and Russian). We allocate one random language to each client, which naturally introduces domain shift among clients.
@@
  • QA: The QA datasets consist of four commonly used question-answering datasets, i.e., SCIQ (Welbl et al., 2017), OpenbookQA (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), ARC-Easy and ARC-Challenge (Bhakthavatsalam et al., 2021). The datasets are transformed into classification tasks where the model determines the correct answer for each question and corresponding choices. We allocate the samples from one random dataset to each client, indicating the domain shift among clients.
B.2 Implementation Details
The CLIP model is pre-trained on the DataComp dataset and is adapted from https://github.com/openai/CLIP; the XLM-Roberta-Base model is adapted from https://huggingface.co/xlm-roberta-base; the LLaMA-2-7B model is adapted from https://huggingface.co/meta-llama/LLaMA-7b-chat-hf. In all training tasks, we freeze the embedding layers of the model and fix the classifier as commonly selected layers (Lee et al., 2019b). The values of adopted hyperparameters are summarized in Table 5. For the proposed method, we tune the value of $\lambda$ from {1, 5, 10, 100, 500, 1000}.